please help me solve this X+Y+Z=22
X+Y=15
3X+Y+Z=37

Answer by Alan3354(69443)

X+Y+Z=22
X+Y=15
3X+Y+Z=37
-------------
X+Y+Z=22
X+Y=15
----------- Subtract
z = 7
-------
3X+Y+Z=37
3x+y+7 = 37
3x+y = 30
x +y = 15
----------- Subtract
2x = 15
x = 15/2
y = 15/2
z = 7